OpenStudy (anonymous):

will reward and medal clearest answer to me....and they have to help w/all 4 questions. 6.) 8(6a+7b)-11(2b+8a)=? 8.) 2(a-b)+3(a-b)-4(a-b)=? 14) 4(11a-9b)-7(6a)? 15.) -3(4a-5b)-(a-b)=? this assignment has answer choices I keep trying to solve them, but my answers don't match up with the answer choices. Could Someone Please Help ASAP...

OpenStudy (anonymous):

can you show us your solution so that we can check where you got mistakes?

OpenStudy (anonymous):

8(6a+7b)-11(2b+8b) 42a+52b-22b+88b 42a+30b+88b (42a+118b)<- but that's not right )': number 6

OpenStudy (debbieg):

6*8 is not = 42 Also your problem above has 8a as the last term but in your solution you changed it to 8b

OpenStudy (anonymous):

could u help me resolv it????

OpenStudy (debbieg):

Fix those 2 mistakes and try again.... :)

OpenStudy (debbieg):

8(6a+7b)-11(2b+8a) =48a + 56b - 22b -88a did you get that far?

OpenStudy (anonymous):

ya

OpenStudy (debbieg):

a-ha.... you used +88a

OpenStudy (anonymous):

ya???/

OpenStudy (debbieg):

But you added. you did 48a + 88a.... Look at it.... =48a + 56b - 22b -88a =48a - 88a + 56b - 22b

OpenStudy (debbieg):

But you said you got 136a, so you ADDED 48a + 88a, that's your error.

OpenStudy (anonymous):

26a-32b answer?

OpenStudy (debbieg):

no...... your b term was right before, don't change that! Just fix the a term.... =48a - 88a + 56b - 22b =a(48 - 88) + b(56 - 22) You will have a negative coefficient for the a, because 88>48

OpenStudy (anonymous):

-26a+34b

OpenStudy (debbieg):

how did you get -26? What is 48 - 88?

OpenStudy (anonymous):

-40a+34b

OpenStudy (debbieg):

Remember how to subtract a bigger number from a smaller one: Switch the order, but then make the answer negative. E.g. 48 - 88 = -(88 - 48)

OpenStudy (debbieg):

There you go! That's it.

OpenStudy (anonymous):

mind helping with 5 problems instead of 4 i didn't see one if u don't care helping w/5????????

OpenStudy (debbieg):

I have to go, but if you repost your others one at a time (I think you are supposed to post them separately) I'm sure someone else will help you. Just try them yourself first, and then ask someone to check your work as needed!

OpenStudy (anonymous):